OpenStudy (anonymous):

Having survived the meteor impact, thanks to some last minute evasive maneuvers, the Mathonauts now set their sights on their Interstellar Headquarters. The Interstellar Headquarters orbits the Earth based on the equation y2 + x2 = 40,000. Using the original trajectory of the ship and complete sentences, explain to the pilot how to find where the ship’s path will cross the Interstellar Headquarters’s path. the original equation is y=6x+2

OpenStudy (anonymous):

@Kfins99

OpenStudy (anonymous):

you have two equations with y and x you have to use one and plug it into the other to solve the system. since the second equation is in terms of y you should plug this into the other to solve for x

OpenStudy (anonymous):

thank you @PhysicsGuru but how would the equation look

OpenStudy (anonymous):

instead of \[y^2 + x^2 = 40,000 \] you will plug in 6x+2 for y

OpenStudy (anonymous):

try to do some of the steps... you have to factor the term and solve the quadratic equation for x

OpenStudy (anonymous):

yeah i tried to do that but it doesnt work cuz of the square i get stuck in the process @PhysicsGuru

OpenStudy (anonymous):

ok so we plug in 6x +2 and get \[(6x+2)^2 + x^2 = 40,000\] which can be written as \[(6x+2)(6x+2) + x^2 = 40,000\]

OpenStudy (anonymous):

You have to use FOIL for the term on the left

OpenStudy (anonymous):

but what about the x^2?

OpenStudy (anonymous):

@PhysicsGuru

OpenStudy (anonymous):

after everything is on the left side of the equation you have x^2 left

OpenStudy (anonymous):

After you foil the term on the left you will add the x^2 afterwards ... always go from more complex to least complex when solving for a variable

OpenStudy (anonymous):

tell me how you foil the term on the left

OpenStudy (anonymous):

36x^2+12x+12x+4=40000 am i doing it right? @PhysicsGuru

OpenStudy (anonymous):

Yes however you are missing the extra x^2 term from the left side of the equation

OpenStudy (anonymous):

oh yah so 37^2? @PhysicsGuru

OpenStudy (anonymous):

@satellite73 do you think you can help

OpenStudy (anonymous):

@ns_69

OpenStudy (anonymous):

@goformit100 can you pls helpppp

OpenStudy (goformit100):

You have been helped by so many users. Sir How can I hep you ?

OpenStudy (anonymous):

i gett lost in the process do you get the problem? i just need you to tell me if im doing it right or not.

OpenStudy (anonymous):

@goformit100

OpenStudy (anonymous):

. You can substitute your original trajectory (y=5x+1) into the equation that the Interstellar Headquarters is based on (y2+x2=40,000). You will be with the equation 2(5x+1)+x2=40,000 and you will solve for X. After solving for X, you will have X=3333.12 for your value. You will now take this value back into one of your original equations. You will now have the equation y=5(3333.17)+1 and will solve for Y. Your value for Y will be Y=16666.85. You now know the ship’s path will cross the Interstellar Headquarters’ at the coordinate (3333.17, 16666.85).

OpenStudy (anonymous):

that doesnt work cuz its to the power and i saw this oh yahoo answers also

OpenStudy (anonymous):

ohk that was from yahoo ans fine!

OpenStudy (anonymous):

@phi can plzzzzzzz help

OpenStudy (phi):

\[ y^2 + x^2 = 40,000 \\y=6x+2 \] as already posted up above, replace y in the first equation with 6x+2 \[ (6x+2)^2 + x^2 = 40000 \] can you multiply out (6x+2)(6x+2) ? Did you try ?

OpenStudy (anonymous):

yeah i got 37^2+24x+4=40,000

OpenStudy (phi):

I assume you mean (putting in the x^2) \[ 37x^2+24x+4=40,000 \] subtract 40000 from both sides \[ 37x^2 +24x -39996=0 \] that is a bit ugly looking. But we can use the quadratic formula (with a calculator) to find 2 x values that will work. then you can find the corresponding y values using y= 6x+2 Do you know how to use the quadratic formula?

OpenStudy (anonymous):

x=-B+-SQRT b^2-4ac/2a you want me to solve it and tell you?

OpenStudy (phi):

do it step by step. first what is b^2 - 4ac and the sqr of b^2 - 4ac

OpenStudy (anonymous):

ok 24^2-4(37)(-3996)

OpenStudy (phi):

and we need a number

OpenStudy (anonymous):

ok that = -590832?

OpenStudy (phi):

no, be careful of the signs. minus times minus is plus

OpenStudy (anonymous):

ok so 590832?

OpenStudy (phi):

is your calculator broken ?

OpenStudy (anonymous):

lol sorry

OpenStudy (anonymous):

wait is it 768.655

OpenStudy (phi):

type 24^2-4(37)(-3996)= into google. (But I would try to get the same answer using the calculator... you do want to be able to use a calculator... it is a useful thing to know)

OpenStudy (anonymous):

I feel like your going to kill me if i dont get this one i swear i typed it as you said but the number is still weird here you go 591984

OpenStudy (phi):

oops, it should be 24^2-4(37)(-39996)=

OpenStudy (anonymous):

ok its 591984

OpenStudy (anonymous):

and google says 5919984

OpenStudy (phi):

that is missing a number. you should get 5,919,984

OpenStudy (anonymous):

so thats x?

OpenStudy (phi):

remember the last number is -39,996 NOT -3,996

OpenStudy (phi):

so far you found b^2 -4ac = 5,919,984 you now need the square root of that number (and save it because you will need to use a few times)

OpenStudy (anonymous):

the sqrt is 2433.101724

OpenStudy (phi):

ok , now you know \( \sqrt{b^2 -4ac}= 2433.101724 \) now let's find one of the x values x= (-b + 2433.101724)/(2a)

OpenStudy (anonymous):

lol ok jst hold up im going to solve this

OpenStudy (anonymous):

ok i got 32.5554287

OpenStudy (phi):

ok,now find y using y = 6x+2

OpenStudy (anonymous):

197.3325722

OpenStudy (phi):

yes, so you have one point where the line intersects the ellipse (32.555, 197.333) (rounded to the nearest thousandth) next, find the other x value, using the minus sqrt(b^2-4ac): x= (-b - 2433.101724)/(2a)

OpenStudy (anonymous):

other square you mean x^2 or jst x

OpenStudy (anonymous):

(33 ,197)

OpenStudy (phi):

you lost some minus signs

OpenStudy (anonymous):

what minus signs

OpenStudy (phi):

x= (-b - 2433.101724)/(2a) that will give a negative x. remember the quadratic equation gives two answers because we do ± sqr(b^2-4ac) we are now doing \[ x = (-b - \sqrt{b^2-4ac})/(2a) \]

OpenStudy (anonymous):

oh yeah i totaly forgotot so itss (33,197) (-33,-197)

OpenStudy (phi):

yes. I would round to about 3 decimals, just to show the answer is not an exact integer. (unless the question says round to the nearest integer)

OpenStudy (anonymous):

ok Thank you Soooo muchhhhh Phi! i have been stuck on this for quite a while!! :)

OpenStudy (phi):

Here is a graph of your problem. (I used geogebra, which you can download for free)

OpenStudy (anonymous):

Thank YOU!!!!

OpenStudy (phi):

yw